3. Find the coordinates of the vertex, the axis of symmetry, the intervals the graph is increasing and decreasing respectivelPlease TYPE , step by step
Answer #1

(a) / 2x+315 12x+31252 Isinaning both side we get / = (2x + 3)2 = 25 - 4x² + 12 x + 9 = 25 { using (ath) ² a ² + 2ab +62] =>=> arra - 2 < x-4 <2 3 - 2+4 < x-4+4 L2+4 → 2 < x < 6 - 1 x-61 <3 - 3 < x-6 L3 - 3 +6 <r < 3 +6 => 3 < x 29 (4) 12x- 311, [12